27 January 2010 - LBC-Blue Active Optics Tests during second half

Operator/Observer: JHill, assisted by DMiller

Summary

The first half of this night is covered in Commissioning.Com20100127.

We did a series of LBC Active Optics Experiments outlined by Andrew. These were LBC-Blue tests to measure the LBCFPIA crosstalk terms between Z4-Z11-Z22.

This log is reconstructed from the Operator's Night Report.

Details

07:00 Start reconfiguration of telescope to LBC-Blue

07:23 Telescope ready to go (almost)

Open

07:27 Open in partly cloudy sky

Alarm in Balls for Latch Pin 0708

07:43 Slew to M81 focus field

edit /home/LBTO/.cshrc for new LBCFPIA path

07:48 dofpia, /blueonly but no pupils.

07:52 Clear Active Optics and Secondary Offload

Oops enclosure rotation was off because I stopped HBS for swing arm motion.

07:55 Try again slewing to M81

07:58 dofpia, but pupils too faint

08:10 slew to Virgo Cluster

08:16 slew to G163m50

08:18 dofpia
Third image is 082107
Messages about xtalk and gain all seem to have the same numbers.

Experiment 0

08:30 Send Z11 +1000 083052
-250 083249
-250 083422
-250 083642

08:37 dofpia
Image 2 has clouds, sent bad focus
Image 4 is 084150

Experiment 1

08:44 Send Z11 -1000 084514
Z11 +1750 084641
Z11 -1500 084827
Z11 +1250 085014
Z11 -1000 085151
Z11 +750 085348
Z11 -500 085502
Z11 +250 085613

08:56 dofpia starting with 085756
Second image has clouds
Third image has clouds

09:03 dofpia starting with 090408

Experiment 2

09:06 Z22 -400 090633
Z22 +700 090819
Z22 +700 090819
Z22 -600 091010
Z22 +500 091126
Z22 -400 091343
Z22 +300 091530
Z22 -200 091645
Z22 +100 091913

09:20 dofpia starting 092054

Experiment 3

09:24 Send Z11 -1000 then dofpia starting 092533

09:29 Send Z11 +1000 then dofpia starting 093041

09:36 Send Z11 -750 then dofpia starting 093648
clouds and bad seeing messed it up

09:40 dofpia without new corrections starting 094021

09:43 Send Z11 -750 then dofpia starting 094421

09:48 Send Z11 +750 then dofpia starting 094932

09:53 Send Z11 -500 then dofpia starting xxxxxx

09:55 Send Z11 +500 then dofpia starting 095649

10:00 dofpia 100119 (single iteration)

Experiment 4

10:03 Send Z22 -400 the dofpia starting 100415 false termination for apparent bad seeing (or Z22?)

10:07 dofpia apparent bad seeing again, clouds are thickening

10:10 Send Z22 +300 then dofpia starting 101147 (converged as it should have)

10:13 Send Z22 +300 then dofpia starting xxxxxx CLOUDS - only 1 pupil

Close

10:25 Close for clouds

Balls on latch pin 0708 not working
RIGHT SHUTTER PINS ARE NOT TORQUED

10:40 Send Doug up to override limit on LDG Rotator so we can turn it on. We were hoping for three instrument configurations in one night, but the clouds aren't going to cooperate.

-- JohnHill - 16 Mar 2010
Topic revision: r1 - 16 Mar 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback